# Switti: Designing Scale-Wise Transformers for Text-to-Image Synthesis

### v3

@@ 5.4 Ablation study
Finally, we evaluate the effect of our architectural choices on the image generation quality and inference time. The human evaluation results in Table 4 demonstrate that the Switti is not only more sample efficient than its AR alternative, but also exhibits slightly better visual quality with respect to all aspects of human evaluation. Moreover, disabling classifier-free guidance at the last two scales noticeably reduces defect presence in Switti’s synthesized images without affecting other aspects of evaluation, as illustrated in Figure 11. Nevertheless, it should be noted that enabling CFG at the last scales can still be beneficial. For example, in scenarios where prompts contain text to be rendered in a small font size, guidance can improve the spelling.
@@
In terms of sampling efficiency, as we show in Table 5, disabling CFG on the last two scales reduces the latency by nearly $20\%$, whereas transitioning to a non-autoregressive architecture further reduces latency by an additional $11\%$.
@@
Switti (AR) 112 179
139
Switti 95 161
127
Table 5: Latency comparison of architecture and sampling modifications proposed in Switti.
@@ 6 Limitations and future directions
Adapting to higher resolutions.
This work uses a publicly available RQ-VAE that is not designed for $1024{\times}1024$ resolutions or higher. In future work, we expect to see effective RQ-VAE models for higher-resolution sampling.
@@ 7 Conclusion
Appendix A Basic architecture
Figure 12 illustrates the design of our basic transformer architecture described in  Section 3.1. Note that the normalization layers are applied only to the inputs of the attention and FFN blocks.
Figure 12: Transformer block of the basic architecture.
Appendix B Training details in analysis
For the experiments in  Section 3.2, we train more light-weight models, with $d{=}20$ transformer blocks, resulting in approximately $0.7$B parameters. All models are trained in mixed precision BF16/FP32 for $150,000$ iterations using Adam ($\beta_{1}{=}0.9$, $\beta_{2}{=}0.95$) with a learning rate of $41e{-}$ 4, linearly decaying to $51e{-}$ 5. Batch size is $768$. Image resolution is $256{\times}256$. In the normalized RoPE, we use $\theta{=}10,000$ and max size $128$. For these experiments, we disable the conditioning on the cropping parameters.
@@ Appendix B Training details in analysis
Appendix C Additional prompt switching visualizations
@@
Appendix D Effect of supervised fine-tuning
To evaluate to which extent supervised fine-tuning affects Switti, we conduct a human evaluation study. As illustrated in Table 6, SFT largely improves the generation quality in terms of all aspects.
SFT no-SFT 0.610.06 0.660.06 0.720.06 0.720.06
Table 6: Human evaluation of Switti against its no-SFT version.
Appendix E Visual comparison between original and finetuned RQ-VAE
@@ Appendix I Effect of disabling CFG at different scales
Appendix J Exploring self-attention maps
Below, we analyze attention maps across different scales of the pretrained Switti (AR). Specifically, for each pair of scales $s_{i}$ and $s_{j}$ where $i\geq j$, we calculate the average attention score from $s_{i}$ to $s_{j}$ using the formula $\overline{A}_{s_{i},s_{j}}=\frac{1}{|s_{i}|}\sum_{m\in s_{i},n\in s_{j}}A_{mn}$. Here, $A_{mn}$ represents the original self-attention map, and $|s_{i}|$ indicates the number of tokens at scale $s_{i}$. The visualization, averaged over a batch of textual prompts, is in Figure 13. It can be seen that self-attention among image tokens primarily focuses on the current scale and is significantly weaker for preceding scales.
Figure 13: Visualization of self-attention maps at different scales for Switti (AR). The model attends mostly to the current scale.
Appendix K List of prompts used in our figures
@@
2) “Cat as a wizard”
@@
4) “A lizard that looks very much like a man, with developed muscles, leather armor with metal elements, in the hands of a large trident decorated with ancient runes, against the background of a small lake, everything is well drawn in the style of fantasy”
@@
6) “Phoenix woman brown skin asian eyes silver scales, full body, high detail“
7) “Portrait of an alien family from the 1970’s, futuristic clothes, absurd alien helmet, straight line, surreal, strange, absurd, photorealistic, Hasselblad, Kodak, portra 800, 35mm lens, F 2.8, photo studio.”
8) “32 – bit pixelated future Hiphop producer in glowing power street ware, noriyoshi ohrai, in the style of minecraft tomer hanuka.“
@@
1) “A beautiful, pleasant, cute, charming young man in a futuristic suit with elements of white and green, a powerful eco-friendly motif, against the background of a beautiful landscape design with an abundance of green vegetation and water, harmony of color, shape and semantic content, light positive emotional coloring, random angle, high detail, photorealism, highly artistic image.”
2) “A squirrell driving a toy car.”
@@
1) “A young badger delicately sniffing a yellow rose, richly textured oil painting.”
2) “3D,game background,beach,shell, conch,starfish,the sea is far away, cloudy day.”
Figure 14: Visual comparison between original RQ-VAE (left) and with fine-tuned decoder (right).
@@
Figure 16: Impact of CFG at different resolution levels. There are $10$ levels: higher index indicates higher resolution. The guidance scale is $6$.

### v4

@@ 5.4 Ablation study
Finally, we evaluate the effect of our architectural choices on the image generation quality and inference time. The quantitative evaluation in Tables 4 and 5 demonstrates that Switti is not only more sample efficient than its AR alternative, but also exhibits slightly better visual quality with respect to all aspects of human evaluation and automated metrics. Moreover, disabling classifier-free guidance at the last two scales noticeably reduces defect presence in Switti’s images without affecting other aspects of evaluation, as illustrated in Figure 11. Nevertheless, it should be noted that enabling CFG at the last scales can still be beneficial. For example, in scenarios where prompts contain text to be rendered in a small font size, guidance can improve the spelling.
@@
COCO MJHQ
Setup PS $\uparrow$ CLIP $\uparrow$ IR $\uparrow$ FID $\downarrow$ PS $\uparrow$ CLIP $\uparrow$ IR $\uparrow$ FID $\downarrow$ GenEval
Switti (AR) 0.227 0.351 0.91 18.7 0.217 0.375 0.76 8.3 0.62
Switti w/ CFG 0.228 0.353 0.90 17.8 0.218 0.375 0.75 8.1 0.62
Switti 0.229 0.355 0.96 18.2 0.219 0.378 0.81 8.1 0.62
Table 5: Impact of the acceleration modifications at $1024{\times}1024$.
@@
As we show in Table 6, a non-autoregressive architecture improves sampling efficiency, reducing latency by around $21\%$. Disabling CFG on the last two scales further decreases latency by an additional $32\%$. Moreover, by removing the causality in Switti, we free up to $2.3$ GB of GPU memory (in half precision), required to store the KV caches of the self-attention layers during the sampling for a single image.
@@
Switti (AR) 2.3 465 908
1.2 640
Switti 0 369 735
0 540
Table 6: Efficiency comparison of architecture and sampling modifications proposed in Switti.
@@ 7 Conclusion
Appendix A Training details
For all our experiments, we use FSDP with a hybrid strategy for effective multi-host training and mixed precision BF16/FP32. To additionally reduce memory usage and speed up the training steps, we use precomputed textual embeddings from the text encoders. We use AdamW optimizer with ($\beta_{1}{=}0.9$, $\beta_{2}{=}0.95$). In the normalized RoPE, we use $\theta{=}10,000$ and max size $128$.
A.1 Pretraining
Data.
Starting from ${\sim}6$B image-text pairs from the web, we filter out images of low aesthetic quality, based on the AADB [34] and TAD66k [20] aesthetic filters. We additionally consider sufficiently high-resolution images with at least $512$px on each side. The resulting dataset contains central crops of the images with an aspect ratio in $[0.75,1.33]$. The images are recaptioned using the LLaVA-v1.4-13B, LLaVA-v1.6-34B [44], and ShareGPT4V [9] models. The best caption is selected according to OpenCLIP ViT-G/14 [62].
Technical details.
During the first stage of pretraining, we train $d{=}30$ models on $256{\times}256$ resolution for $400$K iterations using a batch size of $2,560$. This stage takes ${\sim}25$K NVIDIA A100 GPU hours. We start with a learning rate of $41e{-}$ 4 with a linear decay to $51e{-}$ 5.
Next, we train the models on $512{\times}512$ resolution for $200$K iterations using a batch size of $768$ and a learning rate of $51e{-}$ 5, linearly decaying to $75e{-}$ 7. This stage takes another ${\sim}12$K NVIDIA A100 GPU hours.
A.2 Ablation experiments
For the experiments in Section 3.2, we train more light-weight models, with $d{=}20$ transformer blocks, resulting in approximately $0.7$B parameters. All models are trained for $150,000$ iterations with a learning rate of $41e{-}$ 4, linearly decaying to $51e{-}$ 5. Batch size is $768$. Image resolution is $256{\times}256$. For these experiments, we disable the conditioning on the cropping parameters.
@@ A.2 Ablation experiments
Appendix B Normalization ablation
In this section, we investigate, how the choice of normalization affects the final model performance. In Table 7, we compare our $d{=}20$ version of Switti (AR), trained on $256{\times}256$ against the same version with switched normalization functions (i.e RMSNorms are used for QK-normalization in attention, and LN for input/output normalization) and against the version without “sandwich” normalization. We find the difference in models’ performance insignificant.
Model ($d{=}20$) PickScore $\uparrow$ CLIP Score $\uparrow$ FID $\downarrow$
Switti (AR) 0.206 0.312 10.8
LN $\leftrightarrow$ RMSNorm 0.206 0.313 11.1
w/o “sandwich” 0.206 0.311 11.0
Table 7: Ablation of normalization choices for $d{=}20$ models trained on $256{\times}256$ images.
Appendix C Training loss analysis
Figure 12: Comparison of training losses at various scales on $512{\times}512$ resolution for $d{=}30$ models.
We find that both models converge to similar total losses at the end of the training, as illustrated in Figure 12. However, Switti (AR) has slightly lower values of loss at later scales, while the non-AR version exhibits marginally better performance at the earlier ones. This pattern holds for all configurations of our model during all stages of the training.
This phenomenon combined with our observation in Section 4.2 that around 80% of image tokens can be replaced without affecting the final image visual quality, and the fact that the final non-causal model slightly outperforms its AR counterpart, indicates that the accuracy at the last scales is less crucial.
Appendix D Additional prompt switching visualizations
@@
Appendix E Visual comparison of the finetuned RQ-VAE
@@ Appendix H Visual comparison against T2I models
Model
Generator
size, B
N steps
1 step,
ms/image
Full,
s/image
Distilled Diffusion Models
SDXL-Turbo 2.6 4 12.4 0.25
DMD2 2.6 4 12.4 0.25
Diffusion Models
SDXL 2.6 25 12.4 0.87
SD3-medium 2.0 28 16.8 0.93
Autoregressive Models
Lumina-mGPT 7.0 1024 224.2
LlamaGen 0.8 1024 3.82
HART 0.7 10 4.7* 0.06
Switti (AR) 2.5 10 11.2* 0.14
Switti 2.5 10 9.5* 0.13
*Time averaged over 10 steps.
Table 8: Comparison of $512{\times}512$ image generation times.
@@ Appendix I Effect of disabling CFG at different scales
Appendix J List of prompts used in our figures
@@
2) “A lizard that looks very much like a man, with developed muscles, leather armor with metal elements, in the hands of a large trident decorated with ancient runes, against the background of a small lake, everything is well drawn in the style of fantasy”
@@
4) “Cat as a wizard”
@@
6) “a human face wearing a massive aztec headdress. With hyper intricate war paint on it’s face, detailed texture, vibrant colors, solid background, hyper realistic, 8K“
7) “32 – bit pixelated future Hiphop producer in glowing power street ware, noriyoshi ohrai, in the style of minecraft tomer hanuka.“
8) “Portrait of an alien family from the 1970’s, futuristic clothes, absurd alien helmet, straight line, surreal, strange, absurd, photorealistic, Hasselblad, Kodak, portra 800, 35mm lens, F 2.8, photo studio.”
@@
1) “the girl is 36 years old. hairstyle square, blonde hair, slightly curly. light blue jeans. the top is a classic mint-colored jacket. a brooch in the form of a bird is attached to the jacket. there are silver earrings in her ears. a small leather shoulder bag in dark brown color.”
2) “a squirrel and an acorn”
@@
1) “a drawing of a house on a mountain”
2) “a wooden jewelry box and a fabric rug”
Figure 13: Visual comparison of $512{\times}512$ image reconstruction between an original RQ-VAE (left) and the one with a fine-tuned decoder (right).
Figure 14: Visual comparison of $1024{\times}1024$ image reconstruction between an original RQ-VAE (left) and the one with a fine-tuned decoder (right).
@@
Figure 16: Impact of CFG at different resolutions. There are $14$ scales: higher index indicates higher resolution. The guidance scale is $6$.